if I have 70 candies and 2 are mm,what percentage is this

Answer by jim_thompson5910(35256)

2/70 = 1/35 = 0.02857

0.02857 = 2.857% (move the decimal 2 spots to the right)

So roughly 2.857% of the candies are M&M's